OpenStudy (anonymous):

Multiply: 2 the square root of 11 • 8 the square root of 22

zepdrix (zepdrix):

\[\large 2 \sqrt{11}\cdot 8\sqrt{\color{royalblue}{22}} \qquad=\qquad 2 \sqrt{11}\cdot 8\sqrt{\color{royalblue}{11}}\cdot\sqrt{\color{royalblue}{2}}\]Multiplication is `commutative`, meaning, we can multiply these things in any order. So let's rearrange them a little bit.\[\large 2\cdot8\sqrt2\sqrt{11}\sqrt{11}\]Root 11 squared gives us back 11.\[\large 2\cdot8\sqrt{2}\cdot11\]Do you understand how to do the rest of the multiplication?

OpenStudy (anonymous):

What I don't get what you just type???????